# SemSteDiff: Generative Diffusion Model-based Coverless Semantic Steganography Communication

### v1

@@ V-A1 Datasets
We utilize Stego260 datasets[24] to verify the performance of SemSteDiff. Stego260 consists of 260 natural images collected from two public datasets [^1] https://github.com/aisegmentcn/matting_human_datasets [^2] https://www.kaggle.com/datasets/iamsouravbanerjee/animal-image-dataset-90-different-animals and web sources. The images are resized as 512*512.
@@ V-A2 Parameter Settings
In SemSteDiff, the key generation module and coverless steganography module are pre-trained. For the key generation module, private key generation is employed BLIP-1 base model [^3] https://huggingface.co/Salesforce/blip-image-captioning-base to extract descriptions from secret images, and public key generation is employed application programming interface of ChatGPT to dynamically generate public keys. For the coverless steganography module, We use the Stable Diffusion v1.5444https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 in the forward and reverse diffusion process, with the 50 DDIM steps.
@@ V-A3 Validation Framework
The SemSteDiff is verified under the AWGN channel ranged from 0 dB to 10 dB SNR, and three different JSCC frameworks including DeepJSCC[30], NTSCC[42], and SwinJSCC[43] as the semantic codec module. For DeepJSCC and SwinJSCC, the compression ratio is fixed at $1/12$, while for NTSCC the maximum code rate $R$ is constrained to $1/16$ to enable a fair comparison across schemes[42]. Each is trained under Signal-to-Noise Ratio (SNR) values of 0, 22, 44, 66, 88, and $10$ dB, respectively. The training of JSCC models and text of SemSteDiff is performed on NVIDIA RTX 6000 Ada Generation GPU under PyTorch 2.7.1 with CUDA 11.8 enabled.
@@ V-B Complexity Analysis
We analyze the computational complexity of SemSteDiff based on models used in the simulation. The complexity focuses on three main parts: the BLIP-1 base model for private key extraction, the LLM for public key generation, the Stable Diffusion v1.5 for coverless image steganography. The overall complexity is determined by the total employment of these modules.
@@ V-B2 Public key generation
The LLM are responsible for the public key generation. In the simulation, we use application programming interface of ChatGPT to achieve high quality public key response. However, due to the inaccessibility of specific internal model architecture, the exact computational complexity of ChatGPT remains difficult to estimate. To enable practical deployment, future work can consider using lightweight LLMs fine-tuned specifically for key generation, or applying knowledge distillation to extract a task-specific lightweight model from a large-scale LLM. In this part, we abstract the LLM-related computational cost as $\mathcal{O}_{\text{LLM}}$
@@ V-B3 Coverless image steganography
Therefore, the overall complexity of the system is $\mathcal{O}(L\cdot N^{2}\cdot d+L^{\prime}\cdot T^{2}\cdot d+T\cdot C^{2}\cdot H\cdot W)+\mathcal{O}_{\text{LLM}}$.
@@ V-D1 Performance of the proposed scheme
As shown in Figure 5, Figure 6, and Figure 7, the proposed SemSteDiff framework consistently achieves robust performance in both distortion (PSNR and MSE) and perception (SSIM and LPIPS). In terms of distortion, PSNR exceeds 20 dB across most testing SNR values, with a peak of 21.94 dB with SwinJSCC-based SemSteDiff scheme under $\text{SNR}=10~\text{dB}$. Even when integrated with the lightweight DeepJSCC framework, SemSteDiff still achieves a PSNR of 21.20 dB. The trends in MSE also illustrate that SemSteDiff maintains a low reconstruction error, with the minimum MSE reaching 0.00773 for SwinJSCC and remaining below 0.0095 for DeepJSCC and NTSCC in most cases. These results demonstrate the effectiveness of the proposed scheme, which achieves low-distortion recovery even in the presence of poor channel environment. In terms of perception, SSIM scores consistently exceed 0.70, reaching up to 0.7232 in optimal configurations. Moreover, these perceptual similarities remain stable even under mismatching training and test SNRs. Also, LPIPS values remain uniformly low, with the best result of 0.2073 observed under the SwinJSCC framework. These perception results reflect SemSteDiff can preserve most semantic features, which is depended on the strong generation ability of diffusion model. Meanwhile, the analysis across different frameworks confirms that SemSteDiff does not depended on any specific JSCC architecture, but rather serves as a plug-and-play module in semantic transmission layer, which reflect SemSteDiff’s broad applicability and stability for SemCom.
@@
Figure 8 illustrates the comparison of whether employ our scheme to SemCom under $\text{SNR}=10~\text{dB}$. As expected, SemCom models without steganography present achieve notably higher PSNR and lower MSE. For instance, SwinJSCC without SemSteDiff attains peak PSNR of 36.32 dB and the MSE reaches 0.00033, while its counterpart with SemSteDiff achieves PSNR of 21.94 dB and MSE of 0.00773. Similar trends also appear in JSCC and NTSCC frameworks. This conisderable gap in distortion metrics, however, does not imply a failure in reconstruction. It highlights a shift in the optimization target: SemSteDiff sacrifices pixel-level precision to achieve the target of security, enhancing the importance of semantic recovery and perceptual coherence. From this perspective, SSIM offers a more relevant measure of success. As we could find all SemSteDiff employments maintain SSIM values above 0.66 across the SNR range, indicating strong structural consistency. For instance, SwinJSCC-based SemSteDiff reaches 0.7232 at 10 dB, and even the DeepJSCC-based SemSteDiff achieves 0.6662, demonstrating the system’s ability to retain layout and spatial relationships to human perceptions. LPIPS further confirms this trend for all simulations confirms LPIPS scores below 0.28. Moreover, SwinJSCC-based SemSteDiff achieves a minimum LPIPS of 0.2073, highlighting its alignment with human perception even under imperfect reconstruction conditions.
@@
Figure 11: Visualization of image steganography and reconstruction of legitimate receiver and eavesdroppers. The first column shows the secrct image. The second column shows the stego image. The following four columns show the reconstructed images of legitimate receiver and eavesdroppers, respectively.
@@ V-D2 Security evaluation under eavesdropping scenarios
Figure 10 shows the quantified result of SemSteDiff with legitimate receiver and eavesdroppers for SNR = 8 dB. The legitimate receiver performs much better than eavesdroppers through all metrics, demonstrating strong semantic protection. Eavesdropper 1, which has no access to either key, simply intercepts the reconstructed stego image. It leads to the SSIM around 0.6717 and PSNR around 19.11 dB, but LPIPS remains high at 0.3, which shows similar structural of secret image but different semantic meaning. This phenomenon is caused by our key design strategy: the public key and private key are paraphrased to be semantically distinct but visually similar, enabling the diffusion model to generate natural-looking images to confuse the eavesdroppers. Eavesdropper 2 performs significantly worse with PSNR drops to 16.49 dB, SSIM to 0.5969, and LPIPS increases to 0.3777. It only has a correct public key but an incorrect private key, causing the output shifts entirely away from the secret image. Eavesdropper 3 with no private key performance similar in metric as eavesdropper 2. Although the public key provides some semantic information, the absence of private key guidance results in the generated images retaining only the general contours and posture similar to the secret images, the generated content is completely unrelated.
@@
Figure 11 illustrates the visualization of a subset of images reconstruction with legitimate receiver and eavesdroppers in SNR = 8 dB of NTSCC model. The legitimate receiver is able to reconstruct images that preserve both structure and semantics across all categories, including landmarks, animals, natural scenes, and etc. However, the eavesdroppers consistently fail to infer the correct content. Notably, even when visual textures and contours appear similar, the semantic essence is either distorted or entirely incorrect in eavesdropper outputs, demonstrating that the generative process cannot be easily manipulated without access to both semantic keys. Consider the sixth row of Figure 11, the private key is “chimpanzee”; the corresponding public key is ”lion”. The legitimate receiver reconstructs the secret image correctly. In contrast, Eavesdropper 1 just recovers the stego image of a lion. Eavesdropper 2 only has correct public key but an incorrect private key “cabin”. In this case, the generated image shows a content of architecture, completely misaligned with the chimpanzee. Eavesdropper 3 without private key only generates an unrelated image of a person with similar posture as chimpanzee. These results highlight the robustness and confidentiality of our SemSteDiff framework. By ensuring that both public and private keys are indispensable for reconstructing secret images, the system effectively prevents unauthorized access, even in scenarios where partial key information is leaked. Moreover, the visual diversity of the selected examples emphasizes that this security property is not limited to specific visual patterns, but persists across various inputs, further validating the reliability of our approach under diverse threat conditions.

### v2

@@ V-A1 Datasets
We utilize Stego260 datasets[51] to verify the performance of SemSteDiff. Stego260 consists of 260 natural images collected from two public datasets [^1] https://github.com/aisegmentcn/matting_human_datasets [^2] https://www.kaggle.com/datasets/iamsouravbanerjee/animal-image-dataset-90-different-animals and web sources. We also utilize the UniStega [50] as supplementary dataset to verify generalizability of our scheme, which are collected from multiple public datasets, including COCO, AFHQ, and CelebA-HQ, as well as other online sources. The images are resized as 512*512.
@@ V-A2 Parameter Settings
In SemSteDiff, the key generation module and coverless steganography module are pre-trained. For the key generation module, private key generation is employed BLIP-1 base model [^3] https://huggingface.co/Salesforce/blip-image-captioning-base to extract descriptions from secret images, and public key generation is employed application programming interface of ChatGPT 4o to dynamically generate public keys. The prompt is set as “Given the following semantic description extracted from an image, generate a public description that preserves only category-level or scene-level semantics, and replaces instance-specific attributes such as appearance details, spatial relations, and identities with another different attribute” for ChatGPT to generate public keys. For the coverless steganography module, we use the Stable Diffusion v1.5444https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5 in the forward and reverse diffusion process, with the 50 DDIM steps.
@@ V-A3 Validation Framework
The SemSteDiff is verified under the AWGN channel ranged from 0 dB to 10 dB SNR, and three different JSCC frameworks including DeepJSCC[2], NTSCC[8], and SwinJSCC[49] as the semantic codec module. For DeepJSCC and SwinJSCC, the compression ratio is fixed at $1/12$, while for NTSCC the maximum code rate $R$ is constrained to $1/16$ to enable a fair comparison across schemes[8]. Each is trained under Signal-to-Noise Ratio (SNR) values of 0, 22, 44, 66, 88, and $10$ dB, respectively. The training of JSCC models and test of SemSteDiff is performed on NVIDIA RTX 6000 Ada Generation GPU under PyTorch 2.7.1 with CUDA 11.8 enabled.
@@ V-B Complexity Analysis
We analyze the computational complexity of SemSteDiff based on models used in the simulation. The theoretical complexity focuses on three main parts: the BLIP-1 base model for private key extraction, the LLM for public key generation, the Stable Diffusion v1.5 for coverless image steganography. The overall complexity is determined by the total employment of these modules. We also measure the overhead consumption of SemSteDiff including inference latency, peak GPU memory consumption, and floating-point operations (FLOPs) under batch size of 1 and input resolution of 512×512 on a NVIDIA RTX 6000 Ada Generation GPU. The results are shown in Table III.
TABLE III: Overhead consumption of major modules in SemSteDiff, where FLOPs of latent diffusion are calculated per sampling step. As the overhead of ChatGPT is inaccessible, we choose lightweight LLM of ChatGLM [11] to measure the Overhead deployment of generating public key.
Module FLOPs Precision Latency (s) Memory (MB)
Private key extraction module 129.28G FP32 0.10 921.02
FP16 0.09 476.05
Semantic codec module 4.95G FP32 0.0017 25.25
FP16 0.0012 22.17
Coverless steganography module 677.76G FP32 9.98 6227.86
FP16 3.88 3258.50
Public key generation module 0.96 11968.41
@@ V-B2 Public key generation
The LLM is responsible for the public key generation. In the simulation, we use application programming interface of ChatGPT to achieve high quality public key response. However, due to the inaccessibility of specific internal model architecture, the exact computational complexity of ChatGPT remains difficult to estimate. In this part, we model the LLM-based public key generation as an inference process of a decoder-only Transformer with key–value caching enabled. The private key is provided as an input prompt consisting of $T_{\text{in}}$ tokens, and the LLM autoregressively generates a public key with $T_{\text{pub}}$ tokens. The LLM is composed of $L_{m}$ Transformer layers with hidden dimension $d_{m}$. At each decoding step, the dominant computation arises from the linear projections and feed-forward operations within each Transformer layer, as well as the attention interaction between the newly generated token and the cached representations of the input private key. As a result, the overall computational complexity of public key generation is approximated as $\mathcal{O}_{\text{LLM}}\approx\mathcal{O}(L_{m}\cdot T_{\text{pub}}\cdot d_{m}^{2})+\mathcal{O}(L_{m}\cdot T_{\text{pub}}\cdot T_{\text{in}}\cdot d_{m})$. It is noted that the complex LLM-based key generation can be executed offline stage to decrease the computational cost in practical deployment.
@@ V-B3 Coverless image steganography
Therefore, the overall complexity of the system is $\mathcal{O}(L\cdot N^{2}\cdot d+L^{\prime}\cdot T^{2}\cdot d+T\cdot C^{2}\cdot H\cdot W)+\mathcal{O}_{\text{LLM}}$. It is noted that the key generation is not involved in the diffusion sampling and JSCC processes. Instead, it can be advanced employed before transmission to prepare the relative keys separately, which means the BLIP and LLM could operate only once in practical deployment, significantly reducing computational cost during real-time transmission. Therefore, its computational cost does not scale with channel conditions, coding rate, or diffusion steps. Although additional overhead arises from interactions between modules, such overhead does not dominate the overall computational efficiency of the proposed framework.
@@ V-C Performance Metrics
5) CLIP Score: To further evaluate the semantic security of SemSteDiff, we also adopt the CLIP score [32] to measure the semantic consistency between reconstructed images and private keys, which reflects the ability of legitimate receivers and eavesdroppers to recover secret information. CLIP score measures the alignment between an image and a textual semantic description in a shared embedding space, providing a direct indicator of whether meaningful semantic information can be recovered. It measures the cosine similarity between reconstructed image $\hat{\mathbf{x}}$ and private key $K_{priv}$
$\mathrm{CLIP}(K_{priv},\hat{\mathbf{x}})=\cos\left(\mathbf{v}_{priv},\mathbf{v}_{img}\right)=\frac{\mathbf{v}_{priv}^{\top}\mathbf{v}_{img}}{\|\mathbf{v}_{priv}\|_{2}\|\mathbf{v}_{img}\|_{2}},$ (42)
where $\mathbf{v}_{img}$ and $\mathbf{v}_{priv}$ are normalized embedding vectors of reconstructed image $\hat{\mathbf{x}}$ and private key $K_{priv}$, respectively. A higher CLIP score indicates stronger semantic alignment between the reconstructed image and the target semantic description.
@@ V-D1 Performance of the proposed scheme
As shown in Figure 5, Figure 6, and Figure 7, the proposed SemSteDiff framework consistently achieves robust performance in both distortion (PSNR and MSE) and perception (SSIM and LPIPS). In terms of distortion, PSNR exceeds 20 dB across most testing SNR values, with a peak of 21.94 dB with SwinJSCC-based SemSteDiff scheme under $\text{SNR}=10~\text{dB}$. Even when integrated with the lightweight DeepJSCC framework, SemSteDiff still achieves a PSNR of 21.20 dB. The trends in MSE also illustrate that SemSteDiff maintains a low reconstruction error, with the minimum MSE reaching 0.00773 for SwinJSCC and remaining below 0.0095 for DeepJSCC and NTSCC in most cases. These results demonstrate the effectiveness of the proposed scheme, which achieves low-distortion recovery even in the presence of poor channel environment. In terms of perception, SSIM scores consistently exceed 0.70, reaching up to 0.7232 in optimal configurations. Moreover, these perceptual similarities remain stable even under mismatching training and test SNRs. Also, LPIPS values remain uniformly low, with the best result of 0.2073 observed under the SwinJSCC framework. These perception results reflect SemSteDiff can preserve most semantic features, which is depended on the strong generation ability of diffusion model. Meanwhile, the analysis across different frameworks confirms that SemSteDiff is not depended on any specific JSCC architecture, but rather serves as a plug-and-play module. We further evaluate SemSteDiff in an additional dataset UniStega-Content [50], and the experiment results are shown in Table IV and Figure 8. According to Table IV, SemSteDiff demonstrates stable performance under varying channel conditions. As SNR increases from 0 dB to 2 dB, the PSNR improves from 18.06 dB to 18.36 dB and the SSIM increases to 0.50, indicating consistent reconstruction reliability. The visual results in Figure 8 further validate this robustness, showing that the “baseball player” secret image is successfully recovered even at a low SNR of 0 dB, while the stego image effectively maintains the appearance of the ”robot” public key despite severe channel interference. The above discussion reflects SemSteDiff’s broad applicability and generalization for SemCom.
Figure 8: Visualization of SemSteDiff on UniStega, where the private key is ”a baseball player in a white uniform is bent over” and the public key is ”a robot is bending over with his back facing the camera”.
TABLE IV: Performance of SemSteDiff on UniStega, where the semantic codec module is trained at 0 dB in DeepJSCC framework.
SNR (dB) PSNR (dB) SSIM LPIPS MSE
0 18.06 0.48 0.43 0.017
2 18.36 0.50 0.42 0.016
To evaluate the importance of the private key, we simulate the scheme that the coverless SemSteDiff without private key as comparison. As shown in Table V, the PSNR of SemSteDiff is 19.78 dB but the comparison scheme only achieves 16.74 dB at $\text{SNR}=0~\text{dB}$, for private key preserves semantic information to reconstruct the secret image as an guarantee. The comparison scheme is theoretically capable of recovering the secret image based on the invertibility of DDIM, however, it fails to achieve robustness as SemSteDiff under the inference of channel noise and semantic compression. The design of dual semantic key ensures successful employment of coverless steganography in lossy transmission environment, which demonstrates the effective of SemSteDiff.
TABLE V: PSNR comparison of the proposed scheme with and without private key, where the DeepJSCC framework is trained at SNR = 10 dB.
SNR (dB) Without Private Key With Private Key
0 16.74 dB 19.78 dB
2 17.29 dB 20.42 dB
4 17.69 dB 20.68 dB
6 18.00 dB 20.98 dB
8 18.02 dB 21.08 dB
10 18.10 dB 21.20 dB
@@
Figure 9 illustrates the comparison of whether employ our scheme to SemCom under $\text{SNR}=10~\text{dB}$. As expected, SemCom models without steganography present achieve notably higher PSNR and lower MSE. For instance, SwinJSCC without SemSteDiff attains peak PSNR of 36.32 dB and the MSE reaches 0.00033, while its counterpart with SemSteDiff achieves PSNR of 21.94 dB and MSE of 0.00773. Similar trends also appear in JSCC and NTSCC frameworks. This considerable gap in distortion metrics, however, does not imply a failure in reconstruction. It highlights a shift in the optimization target: SemSteDiff sacrifices pixel-level precision to achieve the target of security, enhancing the importance of semantic recovery and perceptual coherence. Under this paradigm, although JSCC coding parameters such as SNR primarily determine the distortion level of the transmitted stego image, the reconstruction quality of SemSteDiff is notably less sensitive to these variations compared to traditional SemCom models. As shown in Figure 9, higher SNR generally preserves more semantic features; however, the diffusion model provides a strong generative prior that partially compensates for channel-induced distortions. This results in a reduced performance gap across different SNRs, indicating the enhanced robustness of SemSteDiff against channel variations at the semantic level. From this perspective, SSIM offers a more relevant measure of success. As we could find all SemSteDiff employments maintain SSIM values above 0.66 across the SNR range, indicating strong structural consistency. For instance, SwinJSCC-based SemSteDiff reaches 0.7232 at 10 dB, and even the DeepJSCC-based SemSteDiff achieves 0.6662, demonstrating the system’s ability to retain layout and spatial relationships to human perceptions. LPIPS further confirms this trend for all simulations confirms LPIPS scores below 0.28. Moreover, SwinJSCC-based SemSteDiff achieves a minimum LPIPS of 0.2073, highlighting its alignment with human perception even under imperfect reconstruction conditions.
@@
Figure 12: Visualization of image steganography and reconstruction of legitimate receiver and eavesdroppers. The first column shows the secret image. The second column shows the stego image. The following four columns show the reconstructed images of legitimate receiver and eavesdroppers, respectively.
@@ V-D2 Security evaluation under eavesdropping scenarios
Figure 11 shows the quantified performance of SemSteDiff with legitimate receiver and eavesdroppers n SNR = 8 dB of NTSCC framework. The legitimate receiver performs much better than eavesdroppers through all metrics, demonstrating strong semantic protection. Eavesdropper 1, which has no access to either key, simply intercepts the reconstructed stego image. It leads to the SSIM around 0.6717 and PSNR around 19.11 dB, but LPIPS remains high at 0.3, which shows similar structural of secret image but different semantic meaning. This phenomenon is caused by our key design strategy: the public key and private key are paraphrased to be semantically distinct but visually similar, enabling the diffusion model to generate natural-looking images to confuse the eavesdroppers. Eavesdropper 2 performs significantly worse with PSNR drops to 16.49 dB, SSIM to 0.5969, and LPIPS increases to 0.3777. It only has a correct public key but an incorrect private key, causing the output shifts entirely away from the secret image. Eavesdropper 3 with no private key performance similar in metric as eavesdropper 2. Although the public key provides some semantic information, the absence of private key guidance results in the generated images retaining only the general contours and posture similar to the secret images, the generated content is completely unrelated. Moreover, the secure performance is also influenced by SNR, which determines the amount of reliable semantic information available in the transmitted stego image. As the test SNR increases from 0 dB to 10 dB, the legitimate receiver’s PSNR improves significantly from 19.98 dB to 21.33 dB, whereas Eavesdropper 1 only sees a marginal rise from 18.76 dB to 19.11 dB. This widening gap reveals that legitimate receivers leverage semantic keys to guide the reverse diffusion process could defend the channel noise, while eavesdroppers lacking such guidance or wrong guidance cannot compensate for this degradation. The influence in JSCC parameters tend to enlarge the semantic gap between legitimate receivers and eavesdroppers, reinforcing the overall security of SemSteDiff.
@@
Figure 12 illustrates the visualization of a subset of images reconstruction with legitimate receiver and eavesdroppers in SNR = 8 dB of NTSCC framework. The legitimate receiver is able to reconstruct images that preserve both structure and semantics across all categories, including landmarks, animals, natural scenes, and etc. However, the eavesdroppers consistently fail to infer the correct content. Notably, even when visual textures and contours appear similar, the semantic essence is either distorted or entirely incorrect in eavesdropper outputs, demonstrating that the generative process cannot be easily manipulated without access to both semantic keys. Consider the sixth row of Figure 12, the private key is “chimpanzee”; the corresponding public key is ”lion”. The legitimate receiver reconstructs the secret image correctly. In contrast, Eavesdropper 1 just recovers the stego image of a lion. Eavesdropper 2 only has correct public key but an incorrect private key “cabin”. In this case, the generated image shows a content of architecture, completely misaligned with the chimpanzee. Eavesdropper 3 without private key only generates an unrelated image of a person with similar posture as chimpanzee. These results highlight the robustness and confidentiality of our SemSteDiff framework. By ensuring that both public and private keys are indispensable for reconstructing secret images, the system effectively prevents unauthorized access, even in scenarios where partial key information is leaked. Moreover, the visual diversity of the selected examples emphasizes that this security property is not limited to specific visual patterns, but persists across various inputs, further validating the reliability of our approach under diverse threat conditions.
TABLE VI: Comparison between legitimate receiver and eavesdroppers on CLIP score, where the semantic codec module is trained at 8 dB in NTSCC framework
SNR (dB) Legit Eaves1 Eaves2 Eaves3
0 28.36 24.83 17.87 18.07
2 28.57 24.87 18.41 18.39
4 28.58 24.82 18.71 19.26
6 28.27 24.75 18.91 19.56
8 28.34 24.63 19.07 19.67
10 28.16 24.68 19.09 19.80
TABLE VII: Steganalysis detection accuracy of XuNet [46].
Methods Detection Accuracy (%)
Baluja[1] 95.12
ISN[25] 56.23
HiNet[16] 55.71
SemSteDiff 52.18
Table VI presents the CLIP-based semantic consistency between the reconstructed images and the secret keys under different SNRs. It can be observed that the legitimate receiver consistently achieves the highest CLIP scores across all SNR values, with scores around 28.2 to 28.6, while all eavesdroppers exhibit significantly lower semantic similarity. These results confirm that SemSteDiff achieves semantic-level security, where meaningful semantic information can be reliably recovered by legitimate users while being effectively concealed from unauthorized eavesdroppers. Furthermore, Table VII shows the steganalysis detection accuracy of representative traditional image steganography methods integrated with the same NTSCC transmission framework. It can be observed that conventional cover-modification-based methods, including Baluja [1], ISN[25], and HiNet[16], exhibit higher detection accuracy under XuNet[46]. In contrast, the proposed SemSteDiff achieves detection accuracy close to random guessing, demonstrating strong resistance against DL-based steganalysis. This comparison verifies that avoiding explicit cover modification could improve covertness of steganography, thereby enhancing the security of semantic communication under eavesdropping threats.
Figure 13: Trade-off between guidance scale $\beta$ and PSNR, where the semantic codec module is trained at 10 dB in NTSCC framework
Figure 13 further investigates the impact of guidance scale $\beta$ in (35) on the security–performance trade-off of SemSteDiff. As $\beta$ increases, the diffusion process is more strongly constrained by the conditioning semantic key, leading to improved resistance against eavesdroppers while maintaining acceptable reconstruction fidelity at the legitimate receiver. However, when $\beta$ exceeds 2, the reconstruction PSNR of the legitimate receiver drops rapidly from 18.13 dB at $\beta=2$ to 12.27 dB at $\beta=5$, whereas the security gain becomes marginal, indicating a diminishing return region where stronger guidance sacrifices communication performance without proportional security improvement. Also, a larger $\beta$ enforces a tighter alignment to semantic keys. The keys only contains general description while limiting semantic detail in diffusion process. The tighter alignment could degrades the availability for preserving specific private details by reducing the entropy of the sampling distribution, which focus the the generative trajectory on a narrow semantic key.